### v2

# Artificial Intelligence HCI Human-Computer Interaction

Abstract.
A clear and well-documented document is presented as an article formatted for publication by ACM in a conference proceedings or journal publication. Based on the “acmart” document class, this article presents and explains many of the common variations, as well as many of the formatting elements an author may use in the preparation of the documentation of their work.
Abstract.
The advent of Artificial Intelligence (AI) tools, such as Large Language Models, has introduced new possibilities for Qualitative Data Analysis (QDA), offering both opportunities and challenges. To help navigate the responsible integration of AI into QDA, we conducted semi-structured interviews with 15 Human-Computer Interaction (HCI) researchers experienced in QDA. While our participants were open to AI support in their QDA workflows, they expressed concerns about data privacy, autonomy, and the quality of AI outputs. In response, we developed a framework that spans from minimal to high AI involvement, providing tangible scenarios for integrating AI into QDA practices while addressing researchers’ needs and concerns. Aligned with real-life QDA workflows, we identify potential for AI tools in areas such as data pre-processing, researcher onboarding, or conflict mediation. Our framework aims to provoke further discussion on the development of AI-supported QDA and to help establish community standards for responsible Human-AI collaboration.
Qualitative Data Analysis, Large Language Models, Interview Study
$c$s: Human-centered computing Interactive systems and toolsccs: Human-centered computing Collaborative and social computing design and evaluation methodsccs: Human-centered computing Collaborative interactionccs: Human-centered computing User studiesccs: Human-centered computing HCI theory, concepts and modelsccs: Computing methodologies Artificial intelligence
1. Introduction
Recently, Artificial Intelligence (AI) tools have grown increasingly popular, driven by the rapid rise of Large Language Models (LLMs). These powerful models can generate coherent, contextually relevant text across diverse topics and perform tasks such as classification and summarization (Wei et al., 2022a; Bubeck et al., 2023). Large Language Models have been widely adopted across domains, including workspaces, from programming assistance to educational tutoring, and are increasingly finding their way into research contexts.
One area where these models hold significant potential is Qualitative Data Analysis (QDA). Human-Computer Interaction (HCI) researchers are beginning to explore the usage of LLMs throughout their research processes for tasks such as conducting QDA (De Paoli, 2024), generating synthetic data and substituting human participants (Hämäläinen et al., 2023; Schmidt et al., 2024; Argyle et al., 2023), creative ideation (Shaer et al., 2024; Oppenlaender and Hämäläinen, 2023), literature reviews, and writing (Kapania et al., 2024a). Commercial tools have also started integrating AI features (e. g., ATLAS.ti, [^1] https://atlasti.com/ MAXQDA, [^2] https://www.maxqda.com/products/ai-assist and nVivo’s Lumivero AI Assistant [^3] https://lumivero.com/product/nvivo-ai-assistant/(Morgan, 2023).
The rapid adoption of AI tools begs the question: which parts of the QDA workflow are suitable for AI support, and how do we make these determinations? These questions are particularly pivotal for CSCW, where QDA plays a dual role: as a methodological tool, and as a site of collaboration, negotiation, and interpretive practice. Although AI tools can be useful, they raise dilemmas about researcher autonomy (Feuston and Brubaker, 2021) and the risks of imperfect AI outputs (e. g., LLMs hallucinating wrong information) (Xiao and Wang, 2021). As a community, it is imperative that we critically assess which tasks in the QDA workflow are appropriate or inappropriate for AI integration, and to ensure its responsible use. To better understand how AI tools can, and should, be integrated into researchers’ QDA workflows, we conducted semi-structured interviews with HCI researchers about their QDA workflows and opinions on AI integration. We found that while researchers’ QDA workflows share structural similarities, they differ significantly based on project specifics, highlighting the need for AI tools that are highly customizable. Researchers were generally open to AI assistance when it enhanced their workflow but stressed the importance of ethical considerations, trust in AI outputs, and maintaining control.
Our contributions are as follows. First, we provide an empirical account of researchers’ QDA practices, pain points, and challenges. Second, we present a framework for integrating AI into QDA workflows, grounded in researchers’ perspectives on the individual stages of QDA. Based on our findings we ideate tangible scenarios for the integaration of AI, including AI as a productivity tool (e. g., calculating inter-rater reliability), as a mediator (e. g., facilitating consensus), or with human approval (e. g., suggesting codebook components with validation by a researcher). Finally, we discuss how this framework can guide the responsible deployment of AI tools that truly support researchers’ needs and preserve the social value of collaborative analysis practices.
2. Background
In this section, we lay out the methodological backgrounds of Qualitative Data Analysis (QDA) and trace the evolution of AI-assisted QDA from basic computational methods to the growing role of Large Language Models (LLMs). We conclude by synthesizing our research questions.
2.1. Qualitative Data Analysis in CSCW and HCI
QDA is a fundamental process in qualitative and mixed-method research within CSCW and HCI, providing systematic techniques for interpreting non-numeric data such as interviews, field notes, videos, or documents (Booth et al., 2012; McDonald et al., 2019). Its primary goal is to juxtapose or integrate different sources to identify patterns, themes, and insights that deepen our understanding of the research subject. At its core, QDA typically involves iterative cycles of data extraction, synthesis, and analysis (Booth et al., 2012).
QDA is shaped by various methodological frameworks that may overlap in scientific practice (see reporting of QDA steps in Section 4.1) but are distinct in theoretical foundations, flexibility, and depth of analysis. While Thematic Analysis (Braun and Clarke, 2006) is a theoretical framework that focuses on coding features relevant to the research question, Content Analysis (Wilkinson, 2000) quantifies specific events or codes, enabling a blend with quantitative approaches. Grounded Theory (Glaser and Strauss, 1967) aims at theory generation through systematic coding directly from the data, and Framework Analysis (cf. Barnett-Page and Thomas, 2009) applies existing frameworks, refining them throughout analysis, useful in applied research contexts that aim to build upon and extend existing knowledge. In contrast, Narrative Analysis (Riessman, 2008) emphasizes the structure and meaning of data within its narrative context. Coding strategies can be deductive (codes follow predefined categories), inductive (patterns identified in the data), or hybrid, enabling researchers to choose an approach that aligns with their objectives. The choice of coding strategy is not merely technical but reflects a researcher’s underlying goals and questions, guiding the research direction and ensuring the chosen method supports the intended outcomes.
Beyond methods, QDA in CSCW and HCI is shaped by interdisciplinary diversity. Researchers bring different epistemological traditions and backgrounds (Fiesler et al., 2019; McDonald et al., 2019), which leads to ongoing debates about transparency, rigor, reliability, and measures thereof, such as Inter-Rater Reliability (IRR),  (Fiesler et al., 2019; McDonald et al., 2019). At the same time, most CSCW and HCI projects involve more than one researcher (McDonald et al., 2019), making QDA a fundamentally collaborative and social practice of negotiating, resolving disagreements, and making sense of ambiguity. This collaboration is often asynchronous and distributed, and supported by digital tools and software (Wallace et al., 2017), increasingly incorporating Natural Language Processing (NLP) and LLM methods. Yet, despite QDA’s centrality, there has been no systematic documentation of how CSCW and HCI researchers actually conduct QDA in practice, or how emerging technologies might reshape it. Our work addresses this gap by synthesizing and extending fragmented conversations into a coherent account of real-world QDA workflows.
2.2. From Computational Support to AI-Assisted Analysis
While CSCW and HCI researchers draw on a variety of tools for analysis (Fiesler et al., 2019) computer-assisted approaches have long supported QDA. Early tools applied basic NLP techniques such as word frequency analysis, keyword matching, and pattern-based rules (Stone et al., 1962; Marathe and Toyama, 2018). While these methods introduced automation, they lacked the capacity to capture complex human interpretations. Subsequent work introduced supervised machine learning for text classification, extending manual coding into larger datasets (Rietz and Maedche, 2021). Unsupervised approaches such as topic modeling uncovered latent themes without predefined categories (Lennon et al., 2021), and some studies reported surprising alignment with qualitative methods like grounded theory (Baumer et al., 2017; Muller et al., 2016). Such combinations of QDA and machine learning can be especially useful for mixed-methods studies and for analyzing large-scale data, though they raise concerns about transparency and interpretability – values central to qualitative research. Recent advances in generative AI, particularly LLMs like ChatGPT, offer new opportunities. These models are trained on massive text corpora and can generate, summarize, and classify language with a high degree of contextual awareness (Wei et al., 2022b). They encode vast amounts of diverse knowledge, making them potentially adaptable to various tasks within qualitative research. LLMs have shown remarkable performance on annotation tasks where little or no labeled data is provided (Gilardi et al., 2023; Törnberg, 2023a) and have been proposed as collaborators in human annotation teams (Ziems et al., 2024a). Recent prototypes integrate LLMs at different points across the QDA workflow, from inductive coding to codebook development (Bano et al., 2024; Tai et al., 2024; Xiao et al., 2023; Morgan, 2023; Rasheed et al., 2024; Gauthier and Wallace, 2022). This evolution highlights not only the opportunities for scaling qualitative work but also the risks it creates for preserving interpretive rigor.
2.3. Potentials and Risks of AI in the QDA Workflow
Recent studies show how AI can assist at multiple points in the QDA process. Cuevas et al. (2025) examine the usage of LLMs for conducting user studies and analyzing the collected data. For labeling and annotating, LLM-based systems apply existing codes from a given codebook with notable consistency, though they often miss subtle contextual judgments (Xiao et al., 2023). Lam et al. (2024) develop a LLM-based text analysis tool that derives interpretable concepts from unstructured text, inspired by traditional topic models. For data exploration and inductive coding, LLMs can surface novel patterns and assist in the initial exploration of themes and codes, potentially enabling a deeper understanding. However, while LLMs can replicate many primary themes identified by humans, they often miss more subtle and complex interpretations (Morgan, 2023; Zhang et al., 2023; Torii et al., 2024). Hamilton et al. (2023) observed complementary patterns between human and ChatGPT-generated codes, with each identifying themes that the other did not. Barany et al. (2024) investigated the use of ChatGPT for codebook creation, finding hybrid approaches to be more reliable. They suggest LLMs can contribute to the improvement of codebooks, but stress that human involvement remains crucial. While combining interpretive and positivist methods can cause epistemic tensions and discussions among researchers (Baumer et al., 2017; Muller et al., 2016; Paulus and Marone, 2024), AI also holds potential for mediating discussions and facilitating collaboration by highlighting coding discrepancies, and suggesting resolutions (Gao et al., 2023; Ganji et al., 2018; Drouhard et al., 2017).
While these efforts illustrate potential, they also expose tensions. Concerns persist around bias (Bail, 2024; Törnberg, 2023b; Ziems et al., 2024b; Anis and French, 2023; Owoahene Acheampong and Nyaaba, 2024; Bano et al., 2023), lack of reliability for non-English languages (Heseltine and Clemm von Hohenberg, 2024; Suter and Meckel, 2024), and the risk of treating AI as “culturally agnostic” despite its privileging of certain perspectives (Paulus and Marone, 2024). Such issues are particularly important when engaging Global South contexts, where AI risks reproducing inequalities in representation (Kumar and Karusala, 2021; Wacharamanotham et al., 2020; Rifat et al., 2023; Jang et al., 2023; Okolo et al., 2022). Thus, while AI holds potential to bridge “big” and “small” data approaches and to enhance efficiency, it raises equally urgent questions about reliability, ethics, and researcher agency.
2.4. Designing for Human-AI Collaboration
Integrating AI into QDA is not simply a matter of automation. It requires careful design for human–AI collaboration in interpretive, context-sensitive work (Jiang et al., 2021). Existing tools vary widely in how they support AI involvement, from minimal assistance to active roles in coding and analysis, with different levels of human involvement. Moreover, designing for researchers differing in proficiency in tool usage as well as AI capability (Lubars and Tan, 2019; Feuston and Brubaker, 2021) poses specific challenges, especially in diverse fields such as CSCW and HCI. For example, Gauthier and Wallace (2022) developed a toolkit to facilitate computational thematic analysis, emphasizing the need to design it accessibly for a community with different proficiencies in tool usage as well as AI. Thus, they call for ethical designs to involve friction to remind researchers of the imperfection of computational models and issues concerning ethics and transparency. These challenges are also identified by Kapania et al. (2024b) who found HCI researchers struggling to identify such issues in their projects and needing guidance. Similarly, Lam et al. (2024) suggest tools to be transparent to the researcher and acknowledging their limitations, e. g., by showcasing explicit inclusion criteria and scores. At the same time, designs should avoid creating excessive “unanticipated labor” for researchers (Feuston and Brubaker, 2021).
While prior studies provide valuable insights into the potentials of AI in QDA (Tai et al., 2024; Morgan, 2023; Rasheed et al., 2024; Lam et al., 2024; Gauthier and Wallace, 2022), there is still limited understanding of how AI can be responsibly integrated into researchers’ real-life QDA workflows, respecting their concerns and preferences (Lubars and Tan, 2019; Feuston and Brubaker, 2021). We are thus in need for comprehensive frameworks addressing these issues (Owoahene Acheampong and Nyaaba, 2024; Kapania et al., 2024b). To address these gaps, we study HCI researchers’ QDA workflows and their perspectives on the involvement of AI in qualitative research. Specifically, we investigate the following research questions:
  • RQ1
    How does HCI researchers’ real-life QDA workflow look like?
  • RQ2
    What are HCI researchers’ concerns and conditions for integrating AI-based support into their QDA workflow?
Our overarching goal is to explore how AI can be responsibly integrated into QDA, respecting the workflows, concerns, and conditions of HCI researchers. By systematically mapping QDA practices (RQ1) and surfacing researchers’ conditions for AI integration (RQ2), we contribute a framework for AI support that is aligned with the real-world QDA workflows of HCI researchers. This framework synthesizes fragmented CSCW debates around QDA and provides concrete guidance for developing AI tools that enhance – rather than disrupt – the collaborative and interpretive nature of qualitative research.
3. Research Method
We conducted a two-step study to explore HCI researchers’ perspectives on integrating AI into QDA. The first step involved semi-structured interviews with 15 HCI researchers, representing a range of QDA experience levels. These interviews were designed to uncover pain points, potentials, and conditions for AI support in QDA, focusing on real-life workflows, challenges, and concerns. To promote transparency and rigor, we preregistered our study on the Open Science Framework (OSF) platform.444see https://osf.io/yzkta/?view_only=c4bb80ff8635434db0e7c3166146fa53 In the second step, we used the findings of the interviews to conceptualize ways AI can be integrated into QDA. We developed a framework to illustrate concrete examples of how AI can be integrated into the QDA workflow and contextualize these examples with how they align (or do not align) with participants’ concerns.
3.1. Study Materials
Our study included a pre-questionnaire and a semi-structured interview with a sketching task. The full interview guide is available in Appendix A. Both the complete pre-questionnaire and the interview guide are available in our anonymized OSF repository.
Pre-Questionnaire
The pre-questionnaire consisted of eight questions, capturing participants’ backgrounds and experience with QDA. It covered their research background, current field of research, current position, tools used for QDA, years of experience working with qualitative data, the number of research projects involving QDA, and the data collection and analysis methods they have employed in the past.
Interview Guide
The interview guide was organized into three parts: (a) participants’ real-life QDA workflows (b) pain points and gain points in QDA (c) perspectives on the potential for AI support
First (a), we explored the participants’ overall research process, from data collection to analysis, with a particular focus on coding. We asked them to reflect on their most recent research project, discussing the types of data used, the methods and tools employed, resources accessed, and, if applicable, collaboration with other researchers. To facilitate this reflection, we asked them to sketch their coding process on a piece of paper (sketching task). This visual representation not only helped them articulate their workflow but also served as a reference in subsequent explanations, allowing them to highlight specific areas where pain points arose or where they saw potential for AI support. Then (b), we delved more deeply into the QDA workflow, focusing on what participants found enjoyable or engaging, as well as what they found tedious or challenging. We asked which steps were particularly time-consuming, the difficulties encountered, and the role of tool support, including both its benefits and limitations.
Lastly (c), we explored participants’ expectations around AI-assisted QDA. We inquired about the conditions under which they would be willing to integrate AI into their workflows, which steps they believed could be automated, and which aspects should remain human-driven. Finally, participants were given the opportunity to share any additional thoughts on the topic.
3.2. Recruitment & Participants
We used purposive and snowball sampling methods (Palinkas et al., 2015) to recruit participants who had conducted at least one study involving QDA through our professional networks within the HCI and Usable Security and Privacy (USP) communities. We approached researchers directly and invited them to refer interested colleagues, leveraging peer identification to build a diverse and qualified sample. Our recruitment strategy was designed to capture variation across research domains and methodological traditions relevant to CSCW and HCI. We categorized participants’ primary research field based on their self-description. To capture a broad range of experiences and perspectives, we took care to recruit participants from diverse group affiliations, different study approaches, topic choices, and educational backgrounds. Our goal was to achieve data saturation regarding participants’ workflows, pain points, and gain points.
Sample Description
Table 1 summarizes the participant sample. We interviewed 15 HCI researchers with backgrounds in computer science, cybersecurity, engineering, and social sciences, all firmly embedded in HCI research, with some specifying to focus on USP. Twelve participants were early career researchers (mostly PhD candidates), and three were senior researchers. Eight had worked on 1-3 QDA projects, four on 4-6, and three on more than 7 projects, with qualitative experience ranging from 1 to 9 years. We recruited participants until data saturation was reached. We reached data saturation around interview 12 and conducted three additional interviews to confirm this.
Table 1. Detailed information on our interview participants’ research backgrounds, their QDA experience, and their primary QDA tools and methods. Field categorizations (HCI, USP, Security) are based on participants’ self-description.
No. Research Background Research Field Seniority QDA Experience Primary QDA Tools QDA Methods
P1 CS / Sec. / Eng. HCI, USP ECR 6 y. 1-3 pr. MAXQDA, Excel, Word, Google Docs/Sheets TA, GT
P2 CS / Sec. / Eng. HCI, USP ECR 3 y. 4-6 pr. MAXQDA TA, CA
P3 CS / Sec. / Eng. USP ECR 4 y. 7-9 pr. MAXQDA, nVivo, Atlas.ti, Excel TA
P4 CS / Sec. / Eng. Security ECR 4 y. 1-3 pr. MAXQDA, Excel, Python GT, Inductive Coding
P5 Social Science HCI ECR 2 y. 1-3 pr. nVivo Different kinds of TA
P6 Social Science HCI ECR 3 y. 1-3 pr. MAXQDA, Excel, Word, Google Docs/Sheets GT, Descriptive Analysis
P7 Social Science HCI ECR 1 y. 1-3 pr. MAXQDA CA
P8 CS / Sec. / Eng. HCI ECR 1 y. 1-3 pr. MAXQDA TA
P9 Social Science USP ECR 3 y. 4-6 pr. MAXQDA, Excel TA, CA
P10 CS / Sec. / Eng. USP ECR 2 y. 4-6 pr. MAXQDA TA, GT
P11 CS / Sec. / Eng. HCI, USP Senior 6 y. 7-9 pr. nVivo, Excel, Google Docs TA, Open Coding
P12 CS / Sec. / Eng. USP Senior 9 y. 10+ pr. MAXQDA TA, GT, Inductive Coding
P13 CS / Sec. / Eng. Security ECR 1 y. 1-3 pr. MAXQDA, Excel CA
P14 CS / Sec. / Eng. HCI, USP Senior 3 y. 4-6 pr. Google Docs/Sheets, Python Scripts TA, GT, Open Coding
P15 CS / Sec. / Eng. USP ECR 3 y. 1-3 pr. MAXQDA, Atlas.ti Inductive Coding, Open Coding, TA
Legend. CS = Computer Science, Eng. = Engineering, HCI = Human-Computer Interaction, Sec. = Security, USP = Usable Security and Privacy, ECR = Early Career Researcher, TA = Thematic Analysis, GT = Grounded Theory, y. = year(s), pr. = project(s).
Participants first completed the pre-questionnaire, which included an overview of the study procedure and a consent form. Afterward, we scheduled interviews. The interviews were semi-structured, conducted face-to-face, and lasted approximately 45 minutes. Interviews were conducted in either English ($n=9$) or German ($n=6$), depending on participants’ native language and preference. Two researchers were present: a lead interviewer who conducted the session and took notes, and a second researcher who created detailed protocols, including occasional verbatim quotes. To protect participant privacy and follow data minimization principles, we chose not to record audio. The lead interviewer, fluent in both German and English, conducted all interviews for consistency. The second researcher varied among four researchers and was either fluent in English or fluent in both languages as needed. To ensure objectivity, they noted down points that stood out without performing any immediate analysis or interpretation. The researchers then compared, merged, and verified the completeness of their notes, ensuring no conflicts arose. Finally, when necessary, the researchers manually translated their synthesized interview protocols into English for subsequent data analysis.
We piloted our study with three participants to evaluate the procedure and interview flow. As no changes were made, these pilot sessions were included in the final dataset, resulting in 15 participants.
3.3. Ethical Considerations & Data Protection
The study was conducted in accordance with the Menlo Report’s ethical principles (Kenneally and Dittrich, 2012) and approved by our institution’s ethics committee. All participants gave informed consent via the pre-questionnaire and again at the start of the interview. Before the interviews began, the interviewer walked the participants through the consent form again and gave them the opportunity to ask questions and asked for their consent again. To protect the privacy of the participants, they were given pseudonyms after completing the questionnaire and all data was pseudonymized. We stored and managed all data according to national data protection regulations and in accordance with our institution’s policies on confidentiality and data protection.
3.4. Data Analysis
Three authors developed an initial set of deductive codes for tools and resources, data types, and participants’ overall willingness to use AI, based on the pre-questionnaire answer options and interview guide. They each coded three protocols for the coding process, pain points, fun points, and concerns and conditions for using AI within QDA to derive common themes. This process involved closely examining the protocols to identify and conceptualize subcategories. All three coders then created the final codebooks by discussing the identified themes and resolving conflicts in a collaborative session. Two authors then coded the entire dataset, resolving disagreements through discussion until full agreement was reached. To help us later illustrate our findings, the coders included example quotations from the protocols in our codebooks, both paraphrased and verbatim.
Additionally, we analyzed the drawings from the sketching task to derive commonalities and differences in the coding process and extracted participants’ suggestions for concrete AI applications as in vivo codes. Two coders reviewed sketches together with protocols to create a combined metaview of current QDA practices. Each participant’s workflow was redrawn and sequentially mapped into a shared metaview. New elements were added as the emerged, recurring ones were skipped. Closely related processes were later merged to reduce complexity. We differentiated elements taking into account the SeeMe-method developed by Herrmann (Herrmann, 2012). SeeMe is a semi-structured socio-technical modeling technique used for task analysis that involves three visualization elements: (1) roles assigned to individuals or teams, (2) activities that are of a more dynamic characteristic and represent change functions, and (3) entities, defined as objects from the physical world. For entities, we excluded tools and software support from the metaview as they were heavily dependent on the participant’s individual preference, but generally involved in each process element.
4. Results
Here, we present HCI researchers’ QDA workflows, including their pain points (Section 4.1) and report participants’ willingness to use AI support along with concerns and conditions (Section 4.2). In Appendix B, we provide the complete codebooks from our interviews, including descriptions, examples, and quantities for each code. Appendix C shows example workflow sketches.
4.1. RQ1: HCI Researchers Real-Life QDA Workflow
4.1.1. Contexts of Studied QDA Workflows
In the pre-questionnaire, participants reported having analyzed open-ended survey responses ($n=12$), interview transcripts ($n=11$), written documents ($n=8$), and social media posts ($n=5$) as the most common data types. Most participants ($n=9$) reported using more than one tool for qualitative coding, including MAXQDA ($n=12$), Excel ($n=6$), and Google Docs ($n=4$). Participants also indicated experience with multiple methodological frameworks, often adapting them based on the nature of the data and their research context. Thematic Analysis ($n=11$), Grounded Theory ($n=6$), and Content Analysis ($n=4$) were mentioned most frequently.
4.1.2. Uncovering Participants’ Actual Coding Workflow
Figure 1. Detailed overview of participants’ real-world coding workflow, consisting of three stages for codebook creation, codebook refinement, codebook application, and different steps taken in inductive and deductive approaches.:
At a high level, our participants’ coding processes generally followed three main steps: (1) Codebook Creation– developing an initial version of the codebook, (2) Codebook Refinement– updating the codebook through discussion, or by applying it iteratively to new, unseen data, and (3) Codebook Application– applying the codebook to the full dataset, with occasional ongoing adjustments. As illustrated in Figure 1, the majority of our participants followed an inductive approach. Only two created their codebooks based on a priori information with deductive coding. Team structures varied: some participants coded alone, while most had a lead researcher handle codebook creation and refinement, with team input during discussions. In some cases, new researchers applied the final codebook to the full dataset.
Typically, participants began by reading through a subset of the data to develop an overview. Depending on the project, either the entire dataset or a specific subset (often around 10% of the data) was then used to develop an initial codebook. This was done either individually, in parallel by multiple coders who later merged their versions, or collaboratively from the start. In certain instances, this codebook was then applied to the rest of the data, followed by iterative codebook refinements, referred to as a “never-ending process,” as the codebook application progressed and new themes emerged or began to overlap. However, most reported iteratively applying the codebook to new small sets of data and refining the codebook along the way by removing codes and reducing granularity until all team members were satisfied with the result. Finally, the codebook was applied to the rest of the data. This allowed the determination of a final code count.
There was an almost equal split between those who mentioned that the codebook was only applied by one researcher and those who did multiple codings to be able to calculate inter-rater reliability (IRR). In some cases, IRR was used during early codebook development to support refinement; in others, it was calculated only at the end for reporting purposes. One participant described using random checks for verification, rather than full double-coding.
4.1.3. Pains and Gains
Researchers discussed several challenges and difficulties (“pain points”) and rewarding aspects (“fun points”) of their QDA workflows. Understanding both helps identify opportunities for AI support that align with researchers’ values and practices.
Pre-processing
Data often needs to be pre-processed, e. g., transcription of audio recordings. Some participants found this a boring and tedious task, especially when the transcriptions had to be manually corrected or segmented. Several participants expressed particularly enjoying the process of familiarizing themselves with the processed data and finding valuable insights and learning opportunities. This deep engagement was often highlighted as a rewarding aspect of QDA.
Codebook Creation and Refinement
Developing the codebook was often described as time-consuming and presented several challenges:
  1. (1)
    Deciding Focus & Initial Exploration. Deciding the focus of the codes was a key difficulty, as they need to align with the RQs and the specific project context. A few participants noted that poorly defined RQs can lead to incoherent codes. The initial stages of coding were described as particularly challenging: “Going through the first data segments can go wrong in so many ways, and then you have to do it all over again. You learn, but you don’t get a result from that.” (P6)
  2. (2)
    Naming and Refining Codes. Naming and refining codes also requires substantial effort, as codes must be concise and distinguishable: “It can be a complicated process, trying to encapsulate something into a code that also applies to other samples. Going back and refining codes can be very iterative, annoying, and a big challenge.” (P3). This challenge was particularly prevalent among less experienced researchers.
  3. (3)
    Managing the Codebook. Several participants reported difficulties in managing the codebook’s complexity as it grows, having concerns about losing sight of the big picture and ensuring that codes remain coherent. Inferring hierarchies also posed challenges because it can be “painful to put them into high-level themes without missing an important point.” (P1). This pain point was particularly pronounced among participants who handled large datasets or repetitive coding tasks.
  4. (4)
    Restructuring the Codebook. Most participants expressed concerns about restructuring the codebook. They dreaded having to go back and adapt codes, update the codebook, or even start anew if the codebook ended up not fitting the data well: ”where the codebook begins is very different from where it ends.“ (P5)
Despite these struggles, several participants acknowledged that the process of developing a codebook is highly educational. Synthesizing patterns and refining themes was also described as exciting and rewarding, offering a sense of accomplishment.
Codebook Application
Participants described codebook application as one of the most time-consuming tasks. They expressed frustration when adjustments to the codebook were necessary later in the process: “When new codes were identified, the codebook had to be updated, and already-coded data had to be revisited to determine if the new code applied.” (P15) Nonetheless, nearly half of the participants perceived the process of applying the codebook to the data as satisfying.
Collaboration
Most participants reported that while collaboration and iterative discussions are essential to their QDA process, those can be tedious and frustrating, particularly in cases where multiple interpretations are possible or different backgrounds are present, making it difficult to reach consensus. A few participants described onboarding others and initial discussions as cumbersome.
On the positive side, participants highlighted the value of sharing their thoughts with others and thereby learning to better understand the data. These discussions were seen as enriching and helping to refine ideas.
Tools
Throughout the QDA process, participants use a variety of tools, some of which they report having difficulty with. For one participant, managing multiple tools was burdensome, as each tool specialized in different tasks, requiring regular data transfers between them. Issues included cumbersome data transfers, file management, import/export limitations, and collaboration features. In particular, MAXQDA’s complexity overwhelmed some users.
Lack of Structure and Guidance
Participants with limited QDA experience often expressed concerns about the lack of structure and guidance in QDA. They reported feeling uncertain about whether they were conducting the analysis correctly, expressing a need for assistance, such as having a second person review their work. This was linked to the perception that coding is subjective and inherently messy, making it difficult to evaluate whether their observations had objective value or were merely personal insights.
Post-Processing
Participants valued the final stages of analysis, particularly the synthesis of patterns and the formulation of initial results, as one of the most fulfilling parts of the process. This stage allowed them to see the culmination of their efforts and begin to draw conclusions from the data. Additionally, some participants said they enjoyed writing the report and communicating their findings, as well as extracting interesting quotes for the manuscript.
Answering RQ1: Real-life QDA workflows exhibit structural similarities but vary with project specifics
Our study provides a detailed view of how HCI researchers perform and perceive QDA in their everyday practice. Our interviews revealed that while participants’ workflows are similar in structure, they vary highly depending on the project, methodology, data type, degrees of cooperation, and approaches in which codebooks are created. As Figure 1 illustrates, deductive coding is a somewhat straightforward process, whereas inductive coding, more data-driven and interpretative, involves several loops. This is also mirrored in the pain points participants highlighted and has implications for the integration of AI into the QDA workflow, as highly unique use cases also need to be addressed. Similarly, other perceived “pains and gains” during the workflow are often not mutually exclusive: what one researcher considers as enjoyable and vital to the research can be another researcher’s biggest annoyance. For example, less experienced researchers expressed a need for guidance and structure in their workflows, reflecting the importance of mentorship and learning in QDA. This highlights the importance of building AI assistants in a configurable and customizable manner, not as a one-size-fits-all solution, but as systems that can adapt to researchers’ varying needs and levels of expertise.
4.2. RQ2: Participants’ Concerns and Conditions for Integrating AI Support
Almost all participants were open to exploring AI tools in their QDA process. However, most of these participants emphasized that certain conditions should be met or that they would still want to have full control over the QDA process. While a few participants were eager to integrate AI into their workflow, one was strongly opposed to the use of AI, stating that they did not see any useful applications for it in their workflow.
Several participants recognized the potential of AI to automate repetitive tasks – such as data preprocessing, organizing raw data, or annotating thousands of survey responses. Others saw an opportunity for AI to contribute ideas and enrich the analytical process with new views and perspectives. They expressed interest in AI’s ability to identify relationships and suggest novel patterns in the data that might be overlooked, complementing human analysis. However, they also expressed various concerns, which we categorized into four main themes: research ethics, personal involvement and responsibility, quality of AI output, and specific requirements of AI tools. We found no correlation between participants’ level of experience in QDA and their willingness to adopt AI or their concerns regarding its integration. Both novice and experienced researchers expressed openness to exploring AI tools while emphasizing the importance of safeguards and customization to align with their workflows.
4.2.1. Research Ethics
Many participants expressed concerns about the ethical implications and scientific rigor associated with introducing AI into the research process. Specifically, they stressed concerns with privacy, transparency, and confidentiality: participants emphasized the need for AI systems to comply with data protection regulations such as the GDPR, ensuring data storage and usage are transparent and secure, especially when third-party model providers were involved. They wanted clear information on where the data would be stored and how it would be used by the companies providing the AI services. One participant suggested research participants should be informed before data collection if AI will be used to analyze their data. Another participant insisted that any data fed into an AI system must be strictly anonymized, with all personal identifiers removed to protect participant privacy. One participant expressed fear that integrating AI into their research could potentially compromise the confidentiality of their work, risking that AI would expose novel insights or ideas derived from the data. Further, participants were worried about the handling and destination of the data, especially regarding sensitive data and topics, noting that AI systems require special supervision when dealing with sensitive topics, as the involvement of AI might shape the interpretation and dissemination of knowledge in these domains.
4.2.2. Personal Involvement and Responsibility
Participants were reluctant to become overly dependent on AI tools, emphasizing the need to remain directly connected with their data and engaged with the analysis. Even with AI assistance, researchers must retain responsibility for their research outcomes. This sentiment reflects a concern for preserving the integrity and independence of academic research, ensuring that AI serves as a supportive tool rather than replacing the researcher’s role. Moreover, conducting QDA involves learning from the process and improving from experiences. As one researcher explained, “doing the coding has helped me immensely with the analysis, I think it is necessary and I would not want to give that up”. Participants agreed that AI tools should enhance human capabilities rather than degrade them or make the process more cumbersome. One participant expressed concerns about job displacement, cautioning against AI fully replacing human coders.
4.2.3. Quality of AI Output
Participants expressed uncertainty about the capabilities of AI tools, particularly regarding the reliability and trustworthiness of their outputs. A key concern mentioned by several participants was the potential for AI to produce false information, which could compromise the integrity of research. They emphasized that the consequences of such errors could propagate into the final analysis and published findings. Participants were also wary of an AI system’s ability to handle qualitative data, noting that it lacked the necessary experience and background for such tasks: “I see problems with validity when working with AI. […] The researcher has a learning process and will apply categories differently over time, AI does not.” This skepticism led to a reluctance to fully rely on AI for data analysis: “I can’t imagine that there is such an [AI] algorithm and that a human doesn’t have to check it.” Concerns about biases that AI might introduce further fueled this mistrust: “If we add bias to it, we can as well use mine; at least then I know what went wrong, and can account for it in the limitations.
4.2.4. Conditions and Requirements
Given their concerns, participants formulated conditions and requirements for the potential use of AI in QDA, emphasizing tools needed to be transparent, customizable, and supportive of human oversight. To meet privacy considerations, a few participants stressed that AI models should be able to operate in offline mode, ensuring that no data would be uploaded or shared with third parties. Addressing the need for personal involvement and responsibility, several participants emphasized the importance of evaluating the performance and reliability of AI tools before fully integrating them into their workflows. Many participants were particularly concerned about retaining control and agency over the analysis process – with several participants stressing the ability to adjust the level of AI’s influence, preferring a more interactive collaboration with AI rather than a fully automated analysis. A majority of participants also expressed a strong desire for mechanisms to check and verify AI outputs, ensuring the results could be trusted. Several participants further wanted explanations for AI-generated codes: “I would want an explanation for each code, or at least an example from the data to show where the decision was coming from.” Other participants wanted to take random samples of AI outputs and check them or compare them to their own results. Generally, many participants expressed a preference for an interactive approach to AI integration, where they could collaborate with the AI rather than relying on it entirely. Some participants welcomed the possibility to modify AI settings or tune outputs to their needs.
Answering RQ2: HCI researchers are open to AI support if it enhances their workflow while addressing ethical, autonomy, and quality concerns
Participants were generally open to AI in QDA, provided it enhanced rather than replaced their skills. They valued AI for offloading repetitive tasks, enabling them to focus on deeper interpretive work. Others saw an opportunity in AI to enrich the analysis with new views and perspectives. Hence, AI was perceived both as a tool for efficiency and as a catalyst for analysis. Participants expressed several concerns and mentioned conditions that had to be met. They suggested developing AI tools that operate offline to allow for greater control over the data’s use and storage. Quality and trustworthiness of AI outputs were another major concern, especially given AI’s current limitations in dealing with nuanced data and the risk of generating false information. Participants stressed the importance of rigorously testing and validating AI tools before integrating them into their research workflows to ensure their reliability in research contexts. Still, they saw potential in AI tools that are developed in alignment with their needs, with a focus on transparency, control, and ethical compliance. Notably, willingness to use AI did not correlate with participants’ experience level, suggesting these conditions are shared across different levels of expertise.
5. A Framework for AI-based Support that Aligns with HCI Researchers’ Real-Life QDA Workflows
Our results indicate a general openness to exploring the potential of AI support for QDA among HCI researchers. However, this willingness is balanced by significant concerns about research ethics, the quality of AI output, and maintaining control over the research process. Based on our results, we identify different roles AI could play within QDA from recurring themes from our interviews regarding the appropriate use of AI in QDA, researchers’ conditions and concerns, and the stages in the QDA workflow where AI could provide benefits. We contextualize these roles in related work. To guide the development of future AI tools and help researchers decide when and how to integrate AI into their workflows, we propose a framework outlining three levels of AI involvement: (a) minimal AI involvement (b) moderate AI involvement (c) high AI involvement We provide concrete examples of how AI can be integrated at various stages of the QDA workflow, summarized in Figure 2.
Figure 2. Our framework for integrating AI into QDA, depicting 3 stages from less to more AI involvement, visualizing possible roles for AI we deduced from the data for each stage.:
5.1. Minimal AI Involvement: Supporting Human-Centric Workflows
At this level, AI has little to no direct impact on the core QDA workflow and primarily serves as a supportive tool.
Human-Only
At the lowest level, the researcher conducts the entire QDA process independently, without AI assistance. This is particularly relevant when data is highly sensitive, context-dependent, or when there’s a lack of trust in AI’s ability to achieve a nuanced understanding. While one participant was reluctant to use AI altogether, others were open to involving AI. Some researchers express skepticism about the use of AI for QDA, expressing concerns about opacity in AI processes (Khalili, 2023), ethical challenges (Ashwin et al., 2023), and the erosion of human interpretive insight (Hitch, 2024).
AI for Basic Technical Tasks
AI is utilized to perform repetitive, non-interpretive tasks that do not require human judgment, allowing researchers to maintain full control over the interpretive aspects of QDA. As P12 noted, “transcription can be done by AI as long as it complies with our privacy practices”. AI can handle tasks like adjusting data formats, cleaning data, and organizing it for analysis. For instance, tools like Whisper [^5] https://openai.com/index/whisper/ can transcribe audio recordings.
AI as an Intelligent User Interface
AI acts as a virtual assistant within QDA software, improving usability by guiding users through software features and tools. P1 explained, “some tools are difficult to get to know”. Examples of intelligent user interfaces include virtual assistants like Microsoft’s Clippy, providing extensive domain knowledge and maintaining a model of the user.
AI as an Intelligent User Interface [Uncaptioned image] Figure 3. AI provides step-by-step guidance within QDA software, suggesting functionalities based on user behavior and answering questions to support with problems in tool usage. :
AI as a Productivity Tool
In this role, AI enhances productivity by automating evaluative tasks, such as summarizing data, quantifying patterns, or performing basic statistical analyses. P2 highlighted that “calculating IRR in MAXQDA can be a big pain. AI could offer a customizable IRR calculation”, enabling quicker insights. Beyond QDA, AI tools have demonstrated utility across diverse data work practices, including writing, translation, and literature review. Tools such as DeepL [^6] https://www.deepl.com/translator assist with seamless translation, while platforms like Rayyan [^7] https://www.rayyan.ai/ streamline systematic literature review processes, enabling researchers to focus on higher-level analytical tasks.
AI as a Productivity Tool [Uncaptioned image] Figure 4. AI generates summary reports of QDA projects, such as identifying the most or least frequent codes, or calculating IRR, allowing the researcher to review these insights and make informed decisions more efficiently. :
AI as a Training Tool
AI can be used to train aspiring qualitative researchers through interactive, hands-on learning and assist in onboarding by familiarizing new team members with coding workflows and providing personalized training exercises. Our findings suggest that less experienced researchers, who often expressed uncertainty about their coding accuracy, could benefit from AI training tools that provide structured guidance and immediate feedback, addressing gaps in expertise. Palea et al. (Palea et al., 2024) propose a tool that generates AI hints to help students improve their coding by understanding annotation errors.
AI as a Training Tool [Uncaptioned image] Figure 5. AI supports new researchers in learning and practicing QDA through personalized, interactive instructions, real-time feedback, and simulated collaboration exercises to develop critical thinking and improve coding consistency. :
5.2. Moderate AI Involvement: Human-AI Collaboration
At this level, AI works alongside human researchers to support and complement their coding efforts. The product of QDA is still based on human research efforts and revolves around their expertise and interpretation.
AI as a Mediator
AI mediates between multiple coders, helping to reach consensus and facilitate collaboration. Complaining about time-consuming and tiring meetings with collaborators, P4 noted that AI could assist by focusing discussions on relevant parts and highlighting interpretive discrepancies, reducing the time and effort spent in meetings. For instance, Aemonium (Drouhard et al., 2017) highlights ambiguities and inconsistencies between multiple coders by providing visual overviews of assigned codes. CodeWizard (Ganji et al., 2018) identifies coders’ certainty and disagreements to improve consensus. CollabCoder (Gao et al., 2024) provides real-time synchronization, acting as a suggestion provider, mediator, and facilitator throughout the workflow.
AI as a Mediator [Uncaptioned image] Figure 6. AI helps researchers resolve coding conflicts and reach consensus efficiently by analyzing discrepancies and suggesting resolutions. :
AI for Validation
Participants often felt uncertain when conducting QDA and saw potential in using AI for validating their coding, e. g., by comparing their results with AI-identified codes or segments. In this role, AI serves as a validation tool, identifying gaps or inconsistencies, and providing a different perspective. By providing a layer of validation, such a tool can address researchers’ concerns on uncertainty and subjectivity in the coding process. For instance, AI could verify the completeness of coding and highlight passages that may need revisiting. Our findings indicate that this role is particularly valued by less experienced researchers, who expressed a greater need for external validation to build confidence in their analysis, emphasizing the potential for AI tools to support skill development and ensure reliability.
AI for Validation [Uncaptioned image] Figure 7. AI helps identify errors and inconsistencies while keeping researchers actively engaged, e.g., by suggesting potentially missing annotations and providing them for review. :
AI for Refinement & Explanations
AI supports refining the codebook by providing names, descriptions, and restructuring. P1 noted that AI could help organize the codebook once the initial data analysis is completed. P2 added that AI might suggest merging or grouping similar codes to increase clarity. For instance, MAXQDA’s AI features [^8] https://www.maxqda.com/blogpost/ai-coding-of-qualitative-data provide explanations for coded segments, aiding researchers in organizing and refining their codebooks.
AI for Refinement & Explanations [Uncaptioned image] Figure 8. AI provides support by offering concise code definitions and names, e.g., for the final codebook, and suiting quotes for codes. :
AI for Analytics
AI provides data-driven insights to support the human-led coding process, e. g., by giving warnings or identifying patterns. For instance, Annota (Palea et al., 2024) marks passages for potential additional improvements.
AI for Analytics [Uncaptioned image] Figure 9. AI identifies patterns in coding and provides warnings to support decision-making, e.g., when codes overlap or themes could be grouped. :
AI Suggestions
AI suggests codes based on its analysis, which researchers can accept, modify, or reject, maintaining control over the process. Our findings indicate that participants were generally open to receiving AI recommendations for new themes and suggestions on where to apply a code. This openness was particularly evident among participants working with large datasets or repetitive coding tasks. These suggestions keep the researcher in control while allowing AI to enhance the efficiency and breadth of the coding process. For example, Marathe et al. (Marathe and Toyama, 2018) and Rietz et al. (Rietz and Maedche, 2021) developed tools that provide coding suggestions based on simple rules, supporting the extension of the coding to unseen data. Gao et al. (Gao et al., 2024) explored collaborative workflows where LLMs propose new codes and code groups.
AI Suggestions [Uncaptioned image] Figure 10. AI analyzes the data and recommends codes which can be reviewed by the researcher and refined by the AI, preserving human agency. :
AI as a Sparring Partner
AI and human researchers work together throughout the QDA process. AI assists at various stages, such as generating and refining codes, while the human researcher guides the overall process. Participants preferred an interactive or conversation-like AI, with P7 suggesting that AI could act as a “sparring partner” for discussing categories and collaboratively assigning codes. For instance, Chew et al. (Chew et al., 2023) propose a holistic role for LLMs in co-developing and revising codebooks and conducting final coding.
AI as a Sparring Partner [Uncaptioned image] Figure 11. Researchers and AI collaborate interactively throughout the QDA workflow, from codebook creation, over codebook refinement, to codebook application. :
5.3. High AI Involvement: AI-Led Workflows
At this level, AI takes the lead in the QDA process, performing tasks autonomously with varying degrees of human oversight.
AI with Human-in-the-Loop
AI primarily conducts QDA tasks, while human researchers provide input, refine outputs, or make adjustments at key stages. As P8 noted, “The AI suggests a codebook, and I give feedback that the tool uses to make an updated suggestion.” This role allows AI to handle much of the workload while human researchers retain control over decisions. Cohn et al. (Cohn et al., 2024) propose a human-in-the-loop prompt engineering approach to collaborative discourse analysis. Dai et al. (Dai et al., 2023) develop an LLM-in-the-loop framework to develop an initial codebook.
AI with Human-in-the-Loop [Uncaptioned image] Figure 12. AI takes the lead in QDA, e. g., creating a codebook draft (left) or performing initial coding (right), with human oversight throughout. :
AI with Human Approval
AI autonomously performs QDA tasks, but human approval is required before finalizing outputs. This ensures AI contributions are vetted for quality and integrity. P13 suggested, “AI could apply my codebook, and I manually review a random subsample.” Recent studies use LLMs as additional coders that are evaluated in a final step by computing agreement metrics with human coders (Tai et al., 2024; Bano et al., 2024; Xiao et al., 2023). Drapal et al. (Drápal et al., 2023) propose a framework in which an LLM predicts and applies code and themes with intermediate reviews by the researcher.
AI with Human Approval [Uncaptioned image] Figure 13. AI creates a codebook draft (top) or assigns codes (bottom). A researcher approves before decisions are finalized and changes are applied. :
AI with Conditional Autonomy
AI operates independently in defined scenarios and refers more challenging cases to human researchers. P4 noted, “I see AI as tooling for low-hanging fruit, so I can focus on more complex codes.” For example, AQUA (Lennon et al., 2021) suggests applying AI coding for codes where AI-human IRR matches inter-human agreement.
AI with Conditional Autonomy [Uncaptioned image] Figure 14. AI autonomously handles routine coding but alerts researchers when it encounters data requiring nuanced understanding or deviating from the specified criteria. :
Full AI Delegation
At the highest level of AI involvement, AI autonomously manages the entire QDA process without human intervention. OpenAI’s Viable is an AI model that can be used for analysis of unstructured data at scale, for example, in businesses that require regular customer feedback evaluations. Rasheed et al. (Rasheed et al., 2024) propose a multi-agent system that takes data sources, autonomously performs the chosen approach of QDA, and outputs a complete analysis. AI tools for QDA must be highly customizable to adapt to the varied tasks and needs of researchers in qualitative research. One promising avenue for achieving this flexibility is through the concept of agentic AI, which envisions AI systems as autonomous agents capable of performing tasks, making decisions, and interacting with users in a dynamic and customizable manner  (Shavit et al., 2023). While this level of autonomy may be suitable for large-scale data processing, it raises concerns around data validity, interpretive depth, and the loss of human oversight (Jiang et al., 2021). Participants stressed the importance of maintaining agency and control over AI-driven processes.
5.4. Summary: AI Tools for QDA Should Balance Support with Ethical and Practical Considerations
While HCI researchers were open to exploring the potential of AI support for QDA, they emphasized the need for responsible design that respects research ethics, quality, and researcher control. On the basis of our interview results and prior literature, we synthesized a framework of varying AI involvement – ranging from minimal to high – that allows researchers to choose the appropriate level of AI assistance based on their specific needs and project context. This approach ensures that AI enhances productivity and provides valuable insights without undermining the researcher’s agency. By incorporating AI in a way that is transparent, ethically sound, and aligned with established research practices, researchers can responsibly integrate AI into QDA workflows, ensuring it serves as a complement rather than a replacement for human expertise.
6. Discussion
6.1. Synthesizing QDA Practices in CSCW
Our findings extend prior discussions of AI-assisted qualitative research by grounding them in a systematic account of how CSCW and HCI researchers actually conduct QDA in practice.
[…]
While AI-supported QDA offers novel opportunities, a major concern regarding the quality of outputs and inherent biases prevails. This is particularly important when analyzing non-English data (Heseltine and Clemm von Hohenberg, 2024; Suter and Meckel, 2024), engaging minority groups, or conducting research in the Global South, as AI is not “culturally agnostic” (Paulus and Marone, 2024) and often privileges dominant or high-frequency representations over others (Feuston and Brubaker, 2021). Perspectives from the Global South and majority world are often marginalized in HCI and CSCW (Kumar and Karusala, 2021; Sturm et al., 2015; Linxen et al., 2021)and in related disciplines (Hasegawa et al., 2024; Sturm et al., 2015). AI-supported QDA risks amplifying this imbalance by embedding dominant interpretative frameworks and overlooking locally relevant meanings (Kumar and Karusala, 2021), particularly when models are trained on Global North or WEIRD datasets (Linxen et al., 2021). These risks are compounded by ongoing challenges in HCI and CSCW around citational justice (Kumar and Karusala, 2021) and transparency of research practices (Wacharamanotham et al., 2020). Addressing them is essential for building culturally sensitive approaches to AI-supported QDA.
6.2. Putting Our Framework Into Practice
How our framework might be used
Both prior research and our empirical findings highlight the dual need to respect the interpretive, collaborative nature of QDA while exploring how AI might responsibly augment it. Building on these insights, we propose a nuanced framework for human-AI collaboration in QDA, with varying roles for AI across increasing levels of involvement, tailored to the specific needs of researchers performing QDA. We derive tangible scenarios for AI support that go beyond coding to include tasks such as data preprocessing, onboarding new team members, mediating conflicts, and acting as a conversational partner. We present this framework as flexible rather than prescriptive, recognizing that many factors must be considered when integrating AI into qualitative research. Our findings contribute to ongoing CSCW discussions around responsible, human-centered AI design. They offer timely insights into how HCI researchers are beginning to re-evaluate the role of AI in light of rapidly advancing LLM capabilities.
Our framework aims to empower those who want to incorporate AI tools into QDA, but are unsure of how to approach it or where to intervene. A naive approach is to jump towards AI doing the qualitative coding itself, but this was a contentious topic for our participants. Instead, our framework offers numerous ways AI tools can be incorporated in ways that facilitate QDA without compromising research autonomy. Researchers who are curious about incorporating AI tools such as LLMs but are worried about how this may taint the QDA process can strive for minimal AI involvement, such as AI as a training tool. For companies developing QDA software (e. g., nVivo or MAXQDA), many of which are already adding AI tools to their software, our framework hopes to inspire additional insights into the numerous ways AI tools can be developed to support QDA. Determining the appropriate level and role of AI integration requires careful consideration of factors like research goals, data types, and QDA methods. Below, we outline key considerations for integrating AI into QDA.
Determining the appropriate level of AI involvement
When designing AI-assisted QDA workflows, one should consider what AI tools are capable of and what they should do (Feuston and Brubaker, 2021; Lubars and Tan, 2019). For example, large-scale or structured datasets may benefit from AI’s ability to handle pattern recognition, while nuanced or sensitive data might necessitate limited AI involvement to maintain interpretive depth and ethical considerations. Research goals – whether they focus on exploration, theory generation, or confirmation – also dictate how AI can best support QDA. While in early-stage research, AI can support the identification of patterns and themes, in later stages, AI could be useful for consistency checks or validating human-coded data.
Deciding when AI should lead or support
Related work shows that coding performance can depend on various factors of complexity, e. g., the required level of judgment and interpretation  (Ortloff et al., 2023; Kirsten et al., 2024). While higher task complexity is an indicator of reduced AI involvement, routine and repetitive tasks like transcription offer the potential for AI leadership, freeing researchers to focus on more complex interpretive tasks. AI should be employed where its capabilities align with the complexity of the task. For example, when the model shows high certainty in a task, it might take a more leading role, while in tasks involving ambiguity or high interpretive judgment, AI could step back. In addition, researchers’ comfort and familiarity with AI tools should guide their level of involvement (Lubars and Tan, 2019; Gauthier and Wallace, 2022). Recent work (Bhatt and Sargeant, 2024; Gauthier and Wallace, 2022; Lam et al., 2024) discusses how restricting AI outputs or displaying performance disclaimers might aid the effective use of AI.
Identifying stages where human oversight is essential
Critical review points must be established to ensure that human oversight is maintained where necessary. Depending on the level of AI involvement, researchers should determine at which stages oversight is crucial -— such as after each coding pass, document review, or when noteworthy patterns are identified. In cases where AI operates autonomously in a conditional mode, scenarios that require human intervention need to be defined, e. g., under user-specified or performance-dependent conditions.
Deciding on the degree of human control
Researchers need to balance the efficiency of AI with the ability to intervene and customize AI outputs. Researchers should decide how much flexibility they require in modifying AI outputs. For example, they might want to customize AI-generated codes, suggestions, or thematic summaries based on their interpretive insights. AI tools should offer options for manual adjustment and refinement to align with the researcher’s perspective.
AI integration across theoretical QDA frameworks
Our study suggests that researchers’ coding workflows may not strictly adhere to specific coding techniques or theoretical frameworks in everyday practice. Hence, designing AI support that is rigidly tied to a specific QDA approach may be counterproductive for many research projects. Instead, flexibility should be prioritized to accommodate diverse and evolving workflows. However, for research projects that rigorously follow a specific QDA framework, specialized or adjustable AI models could enhance research rigor and alignment with theoretical principles. Testing the feasibility and effectiveness of such AI systems could involve training models with data coded using a particular framework or instructing LLMs to adopt a designated theoretical approach when performing coding tasks. Furthermore, examining how AI-generated codebooks differ when different frameworks are applied may provide researchers with new insights into their data, supporting more informed decisions about the most suitable coding framework for their project. When AI is used as a supporting agent rather than the primary coder, its suggestions must align with the theoretical and methodological context defined by human coders. Misaligned suggestions can disrupt the research process and frustrate researchers by introducing inconsistencies. Ensuring compatibility between AI tools and the methodological requirements of a study is therefore crucial for maintaining coherence and usability in QDA workflows.
Evaluating Ethical Implications and Compliance
Integrating AI into QDA also involves evaluating ethical considerations, such as data privacy and security, to ensure compliance with ethical and legal standards and regulations. Additionally, participants in human subject studies should be informed during the consent process if AI will be used to analyze their data. Further, the harms of engaging with AI outputs need to be discussed, and safeguards should be applied when feeding data into AI systems. Finally, depending on the level of AI autonomy, a normative discussion on violations of intellectual integrity is necessary. For example, it may be necessary to report the level of AI usage in research outputs to maintain transparency and accountability in the research process. Researchers already use AI throughout their research process (e. g. (Kapania et al., 2024a; Argyle et al., 2023; Hämäläinen et al., 2023; Oppenlaender and Hämäläinen, 2023)), and popular QDA software increasingly integrates AI features – calling the scientific community, and specifically the CSCW and HCI community working at the intersection of humans and technology, to engage this debate and establish guidelines and best practices, e.g., on transparency. Kapania et al. (Kapania et al., 2024a) discuss the ethical guidelines for navigating these risks and concerns.
Human-Centered AI (HCAI) shares these ethical concerns and offers frameworks that prioritize human factors throughout the design, development, deployment, and use of AI systems (Usmani et al., 2023). Previous studies on automation highlight risks such as over-reliance and blind trust, often compounded by users’ reduced awareness and understanding of system behavior (“out-of-the-loop” effect (Xu et al., 2023)(Sarter and Woods, 1995; Xu, 2007). HCAI advocates human-in-the-loop AI systems and emphasizes “human-controlled autonomy”, advocating for AI systems that are designed to prioritize human oversight, transparency, and empowerment (Usmani et al., 2023; Xu et al., 2023). This approach aims to avoid excessive autonomy while addressing both the traditional challenges of automation and the novel complexities introduced by AI’s autonomous and adaptive characteristics (Shneiderman, 2020). By focusing on maintaining user control and awareness, HCAI promotes mechanisms for interpretable AI outputs, user feedback loops, and customizable settings to ensure that humans remain engaged and capable of intervention (Xu et al., 2023). These principles are highly relevant to the design of AI-assisted QDA workflows, aligning closely with our participants’ expressed concerns and conditions for AI involvement.
6.3. Limitations & Future Work
Our study faces several limitations, primarily those common to interview-based research, including self-report, recall, and social desirability biases. Further studies could benefit from alternative methods, such as observation, screen recordings, diary studies, and subsequent quantitative approaches to measure findings at scale.
Our sample consisted of participants with varying levels of experience in qualitative coding and different data types, which limited our ability to provide a unified, in-depth account of one specific QDA approach. Still, the diversity of reported QDA practices highlights the complexity of AI integration. Additionally, all participants were from HCI research, and many also specialized in security and privacy, potentially skewing our findings toward those more familiar with (emerging) technologies and computing. We similarly recognize that participants’ related motivational factors (Bandura, 2012), their differential attitudes (Ajzen et al., 2019), even their pre- or misconceptions (Gawronski and Strack, 2012; Hammer, 1996), and incidental knowledge (Ferguson and Bargh, 2024) concerning AI tools, can influence their responses. Nonetheless, our sample provided detailed insights into the potential for adaptive AI tools that can cater to both novice guidance and expert customization.
While our framework outlines numerous ways AI tools could be integrated into QDA, we do not empirically evaluate how these integrations could be implemented or how they impact the overall QDA process. Future research should assess the performance and usability of these AI tools in QDA, exploring how effectively they support researchers and whether they facilitate the analysis in meaningful ways.
We also stress the importance of further research on the effects of AI on human decision-making in QDA. Studies have shown that users tend to over-rely on AI recommendations, even in low-involvement scenarios, highlighting the risk of blind acceptance and overreliance (Buçinca et al., 2021).
7. Conclusion
In this paper, we investigated how Artificial Intelligence (AI), particularly Large Language Models (LLMs), can be integrated into Qualitative Data Analysis (QDA) workflows from the perspective of HCI researchers. Our interview study with 15 researchers experienced in QDA revealed the varied and context-dependent nature of real-life QDA practices. Most participants expressed openness to greater AI involvement in their QDA workflows, indicating a notable shift from attitudes observed in studies conducted prior to the advent of LLMs. They were interested in AI tools that can assist with tasks such as coding, summarizing, or facilitating consensus among researchers, while emphasizing the importance of maintaining human oversight. Ethical concerns, including data privacy and the reliability of AI outputs, were prominent, underscoring the need for AI tools to be transparent, dependable, and aligned with established research standards. In response to these needs, we propose a framework categorizing AI involvement in QDA into three levels: minimal, moderate, and high. This framework allows researchers to select the appropriate level of AI assistance based on their specific project requirements and context, ensuring that AI complements rather than replaces human judgment. In doing so, our work contributes to more reliable and ethical research practices that benefit to the broader academic and societal landscape.
Appendix A Detailed Interview Guide
Below, we provide the detailed interview guide for our study.
Welcome & Study Information
  • Welcome the participant and briefly introduce ourselves.
  • Check the informed consent form.
  • Provide information on the study and data handling:
    • Participation is voluntary and can be discontinued or withdrawn at any time.
    • This is an exploratory interview, which will last no longer than 45 minutes.
    • There is no compensation for participation.
    • We collect some personal information for our scientific research, which will be stored in pseudonymized form.
    • Recorded data is subject to the guidelines of the General Data Protection Regulation (GDPR).
  • Ask the participant if they have any questions.
  • Explain the purpose and goal of this research:
    • The purpose of this study is to understand the qualitative coding practices and challenges of HCI researchers, as well as expectations for computer-assisted qualitative coding.
    • Your participation will help us achieve this research goal.
    • Results may be presented in a scientific publication.
  • Inform the participant about the study procedure:
    • Ask if the participant completed the online survey.
    • Next: Interview session on qualitative research experiences, challenges, and anticipations, divided into three parts.
    • We will take notes, and we will ask you to sketch your thoughts on paper.
    • We would like to collect your paper sketches afterwards.
  • Ask the participant if they have any questions.
  • Emphasize that there is no pressure: Workflows do not have to be textbook standard, and there is no judgment.
Interview Session on Qualitative Research Experiences
  • Q1: Can you give us a high-level overview of your research process from data collection to analysis? Please use this paper sheet to sketch your workflow and explain what you are sketching. You can take your time in sketching and explain afterwards if you prefer.
    • Please recall your last research project and describe how you coded and analyzed the data.
    • What kind of data (data type) did you code?
    • Did you use any additional resources when coding?
    • What tools and methodology did you use?
    • How many people are usually involved in this process? How do you collaborate?
  • Q2: What parts of coding do you find the most interesting? Which parts the most tedious? You can use your paper sketch to talk about individual steps and challenges.
    • Which parts do you personally like the most and why?
    • What are your pain points when it comes to qualitative analysis? What do you dislike about these parts?
        At which stage do you spend the most time?
        Which task/step do you enjoy doing the least and why? What are your major challenges/difficulties when coding?
        To give ideas: segmenting, collaborating, iterating & revising, refining codes, etc.
    • What features do you like and dislike most about the QDA tools you are using?
    • Where do your QDA tools support you?
  • Q3: What are your expectations for AI-assisted QDA? Under which conditions and at which steps are you willing to use AI technologies for QDA?
    • Would you be willing to use AI software that partially automates coding? Under which conditions?
    • Where in your research process would you see such an assistive tool?
    • Are there any parts of the QDA process you wish were automated? Are there parts you think should absolutely not be automated?
    • Looking at emergent technologies (think of ChatGPT, GPT-4), how would you envision these Large Language Models in the research process? (Thoughts, concerns)
    • Optionally: What would such a tool look like for you?
Answering Questions & Farewell
  • Is there anything you would like to add? Do you have any questions?
  • Thank you for your participation!
Appendix B Complete Codebooks from Our Interviews
Table 2. Codebook for Tools and Resources based on our interviews with 15 participants. The codebook includes all codes, along with their descriptions, examples, and the number of participants per code.
Code Description Examples Qty.
Google Docs cloud-based application to collaborate on text-based documents in real-time from any device “Google Docs (nice collaboration features)” 1
Google Drive cloud-based storage service to store, share, and access files from any device “Google Drive which is very convenient for sharing” 1
Google Sheets cloud-based application to collaborate on spreadsheets in real-time from any advice “enter data into a Google Sheet and analyze it” 2
Microsoft Word word-processing software to create, edit, and share text-based documents “Shared Word document: as a shared memo” 2
Microsoft Excel software to organize, analyze, and visualize data using spreadsheets “Codes merged into large Excel table” 9
Text Editor software to create, edit, and manage plain text files using basic formatting and coding features “Text editor: build hierarchies, move codes […], take notes” 1
Memos brief written messages for internal communication to convey information, instructions, or reminders “make sure each code has a memo” 1
Transcription Tools software to convert speech into written text to transcribe audio or video recordings “Interviews are conducted, then transcribed”. 4
NVivo qualitative analysis tool to analyze and visualize data from interviews, surveys, and social media “Get the whole website (download) via nvivo” 5
Post Its small, sticky paper notes to write messages that can be attached to surfaces “I used to do it manually […]. I used sticky notes” 2
Atlas.Ti qualitative analysis tool for complex text, audio, and video data using coding and visualization “atlas.ti and nvivo” 4
Dedoose web-based application for qualitative and mixed-methods research to analyze and visualize data through coding and collaboration “Dedoose for other projects: multiple people can code at the same time” 3
Python versatile, high-level programming language, popular for web development and data analysis “Python script for data processing” 4
Notes App software to create, edit, and organize notes using categorization, search, and cloud synchronization “notes app (to upload to iCloud)” 1
MaxQDA qualitative data analysis software for systematic organization, coding, and interpretation of text, audio, and video data “I used MaxQDA” 12
Miro cloud-based platform for teams to brainstorm, plan, and collaborate in real-time on a shared canvas “Move to Miro: Online Post-It Platform” 2
Shared Pad online collaborative tool to create, edit, and comment on documents in real-time “shared pad for shared notes” 1
Inception open-source platform for collaborative annotation and knowledge extraction from text, enabling users to manage labeled datasets for machine learning “INCEpTION for annotations” 1
API / Crawler interface/program to automatically navigate the web to index content and gather data from websites “Understanding API/Crawler” 1
Table 3. Codebook for Data Types based on our interviews with 15 participants. The codebook includes all codes, along with examples and the number of participants per code.
Code Examples Qty.
Audio Recordings “Transcripts of […] audio recordings” 2
Screenshots “I’ve also done website screenshots” 1
Log Files “software log files to track task performance” 1
Drawing & Illustrations “ I used […] NVIVO for coding pictures/images” 1
Social Media Posts “reddit posts” 3
Written Documents “Mostly written data[…], no transcription required” 2
Interview Transcripts “Interview Study” 8
Open-Ended Survey Responses “Open-ended survey responses, approximately 2000” 1
Think Aloud Protocols “Think-Aloud Protocols” 2
Screen Recordings “screen capture must be analyzed manually” 2
Table 4. Codebook for Fun Points based on our interviews with 15 participants. The codebook includes all codes, along with their descriptions, examples, and the number of participants per code.
Code Description Examples Qty.
Interesting Topic Creating and editing interesting things “Working on transcripts is fun when the topic is interesting” 3
Working With People Interaction with others, e.g., during interviews “working with people, lots of contact” 1
Coding Applying the finalized codebook to the collected data sets “I also really like using the codebook once I have it” 8
Finding Interesting Quotes Extract citations that support the identified themes and the analysis “I enjoy […] sorting out of interesting quotes” 3
Synthesizing Main Themes Synthesize key points of interest “When you […] map [themes] to answers to RQs” 3
Learnings From Collaboration Learning from onboarding, collaboration, and training “Training with Coder 4” 1
Mixing Quantitative & Qualitative Data Use of mixed methods “Bringing together mixed methods quantitative and qualitative data” 1
Forming The Codebook Develop, and edit the codebook “Forming the codebook, when we […] see what results we have” 5
Organizing & Finalizing The Codebook Codebook refinement and finalization “Break it down into lower level codes” 1
Learn & Get Deeper Insights Get deeper understanding of a topic “New experience; […] you get deeper insights into user perspectives” 4
Presenting The Results The presentation of the results, e.g., formulation of the findings “Analyze and *present* the results” 2
Analysis Data collection analysis “Analysis is fun” 4
Discussion Interesting discussions with co-workers “Discussing is fun” 7
Familiarizing & Exploring Data Familiarisation and exploration of collected data “getting familiar with the data” 3
Table 5. Codebook for Pain Points based on our interviews with 15 participants. The codebook includes all codes structured by high-level concepts, along with their descriptions, examples, and the number of participants per code.
Code Description Examples Qty.
Codebook Development
$\hookrightarrow$ Deciding Focus Decide on the focus of the analysis and the important aspects of the data “Decide what to focus on” 3
$\hookrightarrow$ Going Through First Data Browsing the first batch of data to come up with codes and get acquainted “Data viewing, initial data creation” 2
$\hookrightarrow$ Restructuring the Codebook Restructure the entire codebook if something goes wrong “Going back and forth when fixing codes” 8
$\hookrightarrow$ Developing Hierarchies Define code hierarchies and relationships, introduce subcodes, etc. “too many codes, painful to put them into […] themes” 2
$\hookrightarrow$ Finding & Naming Codes Codes should be named in a way that differentiates and clarifies them. “sometimes codes can get long” 5
$\hookrightarrow$ Losing Sight of RQs Easy to forget what the overarching research goals and RQs are “what is the RQ? […] easy to lose sight of that” 1
$\hookrightarrow$ Overdetailing the Codebook Codebook is too detailed, too granular, and contains too many codes “Codebooks become too large and detailed” 3
$\hookrightarrow$ Losing Overview of Codes To lose sight of the mass of codes in the codebook “It is difficult to keep track of the entire […] codebook.” 6
Cleaning Transcriptions Cleanup of transcripts can be tedious and time-consuming “ Cleaning up the transcripts is very tedious” 4
Segmenting Split interviews into shorter pieces for analysis (to code) “If I have very noisy [transcripts], that can be very difficult” 2
Synthesis & Analysis Synthesizing and analyzing the data is challenging. “Synthesize: challenging but also exciting and important step” 1
Collaboration
$\hookrightarrow$ Relying on Others Dependence on others may be a problem “you have to rely on another person, who may fail” 1
$\hookrightarrow$ Initial Discussions The initial discussions about the structure of the codebook “Hard at the beginning” 1
$\hookrightarrow$ Iterative Cycles Repeating cycle of discussion, agreement, and interpretation of codes “discussing code about code with the same discussion” 7
$\hookrightarrow$ On-boarding Others Training other coders can be time-consuming “can be a bit tedious in the beginning, especially the briefing/onboarding.” 1
$\hookrightarrow$ Long Discussions Dialogue and discussion take up a lot of time “two people with strong opinions need to agree” 4
$\hookrightarrow$ Interpretation Discrepancies Discussing how coders interpret the differences “difficulties with others when you understand the categories differently” 9
Codebook Application
$\hookrightarrow$ Tedious Process Assigning codes from the codebook to data segments is tedious “Could also be a bit tedious” 6
$\hookrightarrow$ Making Adjustments Later On To update a code, you have to go back and possibly re-code already coded “Going back and forth when fixing codes” 5
IRR Issues Obtaining a good IRR takes time and is complicated “Discussions, just to get IRR” 5
Time-Consuming Process Coding takes time “If the coding process takes a long time” 3
Lack of Structure and Guidance
$\hookrightarrow$ Need For Assistance Participant wishes assistance with coding “A second pair of eyes helps” 4
$\hookrightarrow$ Messiness Data can be complex and confusing “Social media posts can be very unstructured” 3
$\hookrightarrow$ Subjectivity Interpretation of data can be subjective “Codes can be subjective” 1
$\hookrightarrow$ Uncertainty Interpretation of data can be ambiguous “you never know if you are doing it right” 4
Too Many Tools There is not a single tool that fulfills all tasks “throwing [the data] through a lot of tools” 2
Problems With Tools
$\hookrightarrow$ NVivo: Sharing Files NVivo does not share folders and files seamlessly “Sharing […] is a disaster in nvivo” 1
$\hookrightarrow$ Excel: Adequateness Excel does not quite do justice to many tasks “Excel quickly gets too complex for unstructured data” 3
$\hookrightarrow$ MaxQDA: Import/Export Formats MaxQDA imports and exports data incorrectly “Dislikes: […] export/import” 2
$\hookrightarrow$ Audio Transcription: Low Quality The audio transcriptions are of poor quality and difficult to follow “Audio transcriptions could be more accurate” 1
$\hookrightarrow$ File Management Keeping track of all files and file versions is hard “Manual transfer to Excel spreadsheet was tedious” 1
$\hookrightarrow$ MaxQDA: Initial setup and Learning MaxQDA is complicated and not intuitive to use “Dislikes: Initial setup was difficult” 1
$\hookrightarrow$ MaxQDA: Collaboration It is not possible to work on a file together in MaxQDA “you can’t work on a project in sync” 2
$\hookrightarrow$ MaxQDA: Merging Projects Difficult and incorrect merging of projects in MaxQDA “Merging is pretty unintuitive” 2
$\hookrightarrow$ MaxQDA: Crashes MaxQDA crashes during editing “Crashed often, but did not experience any data loss” 2
$\hookrightarrow$ MaxQDA: IRR Computation MaxQDA does not calculate IRR correctly “Limited tool support for incorrect or unexplained results where appropriate” 2
$\hookrightarrow$ MaxQDA: Too Many/Complex Functions MaxQDA has too many and too complicated functions “MaxQDA: Basic use, not using the full feature set, tedious manual process” 3
Table 6. Codebook for Willingness based on our interviews with 15 participants. The codebook includes all codes, along with their descriptions, examples, and the number of participants per code.
Code Description Examples Qty.
Yes Participant is willing to use AI without conditions “Willingness to use AI support: Yes” 2
Conditional Participant would use AI, but immediately mentions concerns/conditions “Yes, but I do not want it to have full control over my data” 10
No Participant is unwilling to use AI in qualitative coding - 1
Unclear Participant did not state willingness “I want to see how it can complement the research process” 2
Table 7. Codebook for Concerns and Conditions based on our interviews with 15 participants. The codebook includes all codes structured by high-level concepts, along with their descriptions, examples, and the number of participants per code.
Code Description Examples Qty.
Scientific Practices
$\hookrightarrow$ Moral / Ethical Concerns
$\hookrightarrow$ Confidentiality New knowledge should be kept secret “ Confidentiality. Especially when something […] is new” 1
$\hookrightarrow$ Sensitive Topics AI may not deal adequately with sensitive information “depending on the sensitivity” 3
$\hookrightarrow$ Privacy Concerns that AI may not be able to comply with the privacy policy “Concerns: privacy, GDPR compliance” 10
$\hookrightarrow$ Personal Involvement Need to be personally involved and know their own data “Do not make people too dependent on the tools” 10
$\hookrightarrow$ Transparency to Participants Participants should be informed about the use of LLMs “tell participants that we are analyzing […] with LLM?” 4
$\hookrightarrow$ Integrity AI can threaten the integrity and independence of academic research “Academic integrity: response can sound good but not be” 2
$\hookrightarrow$ Organizational
$\hookrightarrow$ Implications for ERB The ERB must be informed of the use of AI “Implications for ERB, what tools will be used?” 1
$\hookrightarrow$ Conference Policies Paper may be rejected because conference does not allow AI-coding “the policies & guidelines of conferences […] around LLMs” 2
Features
$\hookrightarrow$ Compatibility Compatibility with existing tools “broad compatibility” 1
$\hookrightarrow$ Control/Agency
$\hookrightarrow$ Comparison with Human Coding should be compared to coding by a human “If the AI came up with the same results [as a human] multiple times” 6
$\hookrightarrow$ Control Codebook Creation Researcher has full control over codebook content “So I would have a special look over his shoulder” 7
$\hookrightarrow$ Verify Outputs The researcher must verify the data output by the AI “look over 20 percent, similar to the human process” 11
$\hookrightarrow$ Offline Usage Offline operation of AI, data is not uploaded to a cloud server “First pass for anonymization model must run locally” 2
$\hookrightarrow$ Collaborative Assistant Work with AI as a collaborator/interactive assistant “you work together, i would go through suggestions” 8
$\hookrightarrow$ Glitch Free It should be possible to work with the tool without any errors or problems “A glitch-free collaboration feature is really what I want” 1
$\hookrightarrow$ Changing AI Parameters Participant wants to be able to tune the AI “fine-tune things, tinker with different variables” 7
$\hookrightarrow$ Explainability AI has an explanation for its response “Can they provide justification or explanation?” 7
$\hookrightarrow$ Evaluation of AI Performance Knowing about AI limitations “I am an observer of how good they are” 5
Meta Concerns
$\hookrightarrow$ Learning From Coding Humans learn by coding, gaining experience over time. AI does not “Qualitative research is very subjective, […] where does it come from when AI does it?” 2
$\hookrightarrow$ Tuning AI Takes Time Checking and adjusting the AI takes time “[AI] should provide support and not extra work” 4
$\hookrightarrow$ Taking Away Jobs Concerns about AI replacing human workers “workplace at risk” 1
$\hookrightarrow$ Costs of AI tool AI tool may be too expensive, leading to disparities between research institutions “cost of these tools could also be a concern” 1
AI Output Quality
$\hookrightarrow$ Hallucinations AI cannot provide completely accurate data and has produced incorrect results “afraid to let it write a whole background section because of wrong citations” 5
$\hookrightarrow$ Low Quality Low level of confidence in AI coding, how to check it “we give it to the machine, how do we know its […] good?” 4
$\hookrightarrow$ Model Will Not Understand AI does not have the knowledge that humans have to understand “ Qualitative coding is not a straightforward approach” 3
$\hookrightarrow$ No Trust in Full AI Control AI-generated results in doubt “Full agency: do not just trust in the results” 5
$\hookrightarrow$ Subjective Data Is Hard for AI It is difficult for AI to interpret and understand subjective statements “there is no way to rank, e.g., values that are more important than others, at least not for a machine.” 5
$\hookrightarrow$ Different Background We learn from experience, AI doesn’t “don’t think the AI understands what anyone means” 6
$\hookrightarrow$ AI Bias The AI could inadvertently develop biases as it ingests data “Qualitative research is very subjective, […] where does it come from when AI does it?” 5
$\hookrightarrow$ Conversations Unlinke With Collaborators Talking to an AI is not the same as talking to a human, and makes codebook development more difficult “Translating what you give to a human coder into what you give to an algorithm” 3
Appendix C Examples of Participants’ Coding Workflow Drawings
Figure 15. Four examples of coding workflow drawings created by our participants during interviews. The drawings varied considerably in the level of detail they provided. Some included in-depth descriptions of specific actions during each step, while others focused more on visuals, providing details of the process only verbally.: